OpenStudy (anonymous):

The length of a rectangle is 4 times its width and the perimeter is 150 feet. What is the width of the rectangle? A. 15 feet B. 60 feet C. 30 feet D. 75 feet

jagr2713 (jagr2713):

give me one second

OpenStudy (anonymous):

The answer is 15 feet. If length is 4 times width, then 2*(4w+w)=150, or 4w+w=75. therefore 5w=75 and w(width)=15feet.

jagr2713 (jagr2713):

Is it D

jagr2713 (jagr2713):

but i got 71

jagr2713 (jagr2713):

150=2(4)+2w

jagr2713 (jagr2713):

p=2L+2W

OpenStudy (anonymous):

you missed the fact that length is 4 times the width. but you wrote it like length is 4 feet.